OpenStudy (anonymous):

lim sin^2(x)/x x->0 find the limit

OpenStudy (tkhunny):

Have you considered: \(\dfrac{\sin^{2}(x)}{x} = x\cdot\dfrac{\sin^{2}(x)}{x^2}\)?

OpenStudy (anonymous):

ah ok thank you

OpenStudy (tkhunny):

Really? That's all it took? :-)

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

to be totally honest i really suck at simplification but i know what to do after its simplified haha

OpenStudy (zzr0ck3r):

I would not call that simplified, but ...

OpenStudy (tkhunny):

Fair enough. Good work. Have to agree with @zzr0ck3r - more of an UNsimplification! Recommendation: Quit sucking! :-)

OpenStudy (anonymous):

wow ok

OpenStudy (zzr0ck3r):

? why would you say that??? @tkhunny are you friends with @theblondeone5 ? if not that was very rude

OpenStudy (tkhunny):

@zzr0ck3r Why is it rude? @theblondeone5: "...i really suck at simplification..." @tkhunny: "Recommendation: Quit sucking! :-)" I'm not seeing the problem with a friendly retort that is EXACTLY on point. More smilies? What?

OpenStudy (zzr0ck3r):

Woops didn't read that sorry @tkhunny

OpenStudy (tkhunny):

All is forgiven.